### v2

# Nemotron-CrossThink: Scaling Self-Learning beyond Math Reasoning

Abstract
Large Language Models (llms) have shown strong reasoning capabilities, particularly when enhanced through Reinforcement Learning (rl). While prior work has successfully applied rl to mathematical reasoning—where rules and correctness are well-defined—generalizing these methods to broader reasoning domains remains challenging due to limited data, the lack of verifiable reward structures, and diverse task requirements. In this work, we propose Nemotron-CrossThink, a framework that systematically incorporates multi-domain corpora, including both synthetic and real-world question-answer pairs, into rl training to improve generalization across diverse reasoning tasks. Nemotron-CrossThink addresses key challenges by (1) incorporating data from varied sources spanning STEM, humanities, social sciences, etc.; (2) applying structured templates (e.g., multiple-choice and open-ended) to control answer-space complexity; (3) filtering for verifiable answers; and (4) optimizing data blending strategies that utilizes data from multiple sources effectively. Our approach enables scalable and verifiable reward modeling beyond mathematics and demonstrates improved accuracies on both math (math-500: +30.1%, amc23: +27.5%) and non-math reasoning benchmarks (mmlu-pro: +12.8%, gpqa-diamond: +11.3%, agieval: +15.1%, supergpqa: +3.8%). Moreover, Nemotron-CrossThink exhibits significantly improved response efficiency—using 28% fewer tokens for correct answers—highlighting more focused and effective reasoning. Through Nemotron-CrossThink, we demonstrate that integrating multi-domain, multi-format data in rl leads to more accurate, efficient, and generalizable llms.
1 Introduction
Large Language Models (llms) have demonstrated remarkable reasoning capabilities across a wide range of tasks, with Reinforcement Learning (rl) playing a crucial role in refining their deep thinking abilities (Hu et al., 2025; Aggarwal & Welleck, 2025; Luo et al., 2025; DeepSeek-AI, 2025; Qin et al., 2024; Huang et al., 2025; Team, 2025b). Recent advances in rl have been particularly successful in mathematical reasoning and coding, where well-defined rules and verifiable correctness criteria enable effective reward modeling. However, extending these techniques to broader reasoning domains presents significant challenges, such as—including limited training data for rl due to the difficulty of defining verifiable rewards, and ensuring generalization across diverse tasks.
Recent work (Hu et al., 2025; Luo et al., 2025; Cui et al., 2025) has shown a way to diversify rl training corpora by collecting datasets from multiple sources. However, they do not evaluate the relative importance of each source for downstream tasks, nor do they explore optimal data-blending strategies to maximize performance gains. Furthermore, prior research has largely focused on mathematical reasoning, overlooking the impact of incorporating non-math reasoning domains in rl-based learning for generalization in out-of-distribution domains. A major challenge in applying rl to general-purpose reasoning tasks lies in designing a verifiable reward model for diverse answer spaces, as unlike mathematical reasoning—where correctness can be objectively verified—other reasoning tasks lack deterministic solutions. Moreover, reasoning process varies across domains and question types. For instance, mathematical problem-solving follows a rule-based, structured, and symbolic approach (Dehaene, 2011), whereas reasoning in fields such as law, physics, social sciences, and history often relies on narrative structures, contextual knowledge, and heuristic search strategies. Additionally, different question formats require distinct cognitive approaches — open-ended questions demand the generation of novel responses from scratch, while multiple-choice (mcq) questions can often be solved more efficiently by evaluating the given options and selecting the most appropriate answer. Incorporating a diverse range of reasoning domains and question types into rl-based self-learning can enhance the broad reasoning capabilities of llms by exposing them to varied cognitive strategies and knowledge structures.
In this work, we propose Nemotron-CrossThink, a systematic way to incorporate multi-domain corpora for rl training that results in better generalization across a wide variety of tasks. As demonstrated in Figure 2, Nemotron-CrossThink comprises of phases that—(a) curate data from diverse sources, including synthetic data from raw web texts (CommonCrawl) and open-source question-answer pairs, spanning STEM, humanities, law, and social sciences (b) apply templates (mcq/Open-Ended) to limit the answer-space for synthetically generated data (c) filter out samples that are infeasible for verifiable rewards (d) prepare blending recipes to combine different sources of data efficiently and finally (e) employ self-learning with rl to refine reasoning capabilities in diverse domains.
Figure 1: Employing self-learning with multi-domain data, Nemotron-CrossThink outperforms baseline models, including domain-specific training (Only Math) and Open-Reasoner-Zero (orz-7B), achieving consistent gains across all reasoning tasks.
Nemotron-CrossThink demonstrates that integrating multi-domain data with different questions formats for rl significantly enhances reasoning ability of llms across diverse reasoning tasks. Notably, models trained with Nemotron-CrossThink not only achieve higher accuracy but also exhibit dynamic response strategies—generating concise answers for general-purpose questions and more detailed responses for math problems—thereby reducing inference cost while preserving task-specific rigor. In addition, Nemotron-CrossThink addresses the challenge of designing verifiable reward for non-deterministic domains by employing different templates on the curated data to limit the nuances in the answer space diversity. This enables scalable, verifiable reward modeling for general purpose reasoning tasks, ensuring that rl-trained models generalize effectively across diverse benchmarks. Furthermore, Nemotron-CrossThink explores a simple yet effective filtering approach to rank general purpose reasoning data based on complexity and shows that training with harder samples further amplifies the impact of rl across all domains.
In summary, our key contributions are as follows:
  • We introduce Nemotron-CrossThink, a novel framework for incorporating multi-domain corpora into rl training, enhancing the generalization of llms across diverse reasoning tasks with substantial gains across both math (math-500: +30.1%, amc23: +27.5%) and non-math (mmlu-pro: +12.8%, gpqa-diamond: +11.3%, agieval: +15.1%, and supergpqa: +3.8%) benchmarks.
  • We demonstrate that applying question/answer templates to constrain output diversity leads to more stable reward modeling. Specifically, using a unified open-ended question format improves performance by 1.21% on average over mixed-format questions, while short-form answer templates outperform long-form ones by 1.20%.
  • We explore optimal data-blending strategies to balance multi-domain corpora and show that only math data is not enough. Blending multi-domain data boosts average reasoning accuracy by up to 1.61% over math-only training and improves response efficiency by reducing token usage by 28%.
  • We propose a simple yet effective model-driven filtering technique that selects harder samples by removing data solvable by smaller models. This leads to an additional 2.15% average accuracy gain for Qwen-2.5-32B, highlighting the scalability of our approach to larger models.
In this paper, we evaluate Nemotron-CrossThink across three dimensions: (1) the effectiveness of different data blending strategies in self-learning (2) whether the blending impact amplifies by filtering and training with more complex data samples (3) the influence of question and answer templates in the downstream performance. Applying Nemotron-CrossThink on different data blends yields substantial improvement over base model, ranging from 8.55% to 13.36% improvement on average across seven diverse general-purpose reasoning and mathematical benchmarks. The most effective blend—constructed using a 2:1 ratio of general-purpose reasoning to math data—achieves the highest average accuracy, improving over the baseline by 13.36% (Figure 1). This underscores the effectiveness of conducting self-learning with combination of data from multiple reasoning domains to enable broader generalization. Our filtering experiment with Qwen-2.5-32B shows a consistent trend, indicating that larger models can further amplify these gains with more complex samples in the data blend (2.15% average improvement), exceeding the improvements observed in the 7B setting. Additionally, our controlled template studies reveal that data formatting decisions play a critical role in model performance. Overall, these findings illustrate that thoughtful choices in data blending, scaling, formatting, and filtering are critical to the success of reinforcement learning with language models. We hope that Nemotron-CrossThink serves as a practical and extensible framework for leveraging multi-domain data to train more capable, reliable, and generalizable models under the rl paradigm.
2 Nemotron-CrossThink: Scaling Self-Learning Beyond Math
Figure 2: Nemotron-CrossThink. We (a) curate QA pairs from from synthetic (Common Crawl) and open-source datasets, categorized into general-purpose reasoning ($\mathcal{D}_{gpr}$) and mathematical reasoning ($\mathcal{D}_{mr}$); (b) apply structured templates to convert data into multiple-choice (mcq) and open-ended formats, promoting diverse reasoning trajectories; (c) filter out unverifiable or ill-formatted responses; (d) train an RL policy using Group Relative Policy Optimization (grpo). The final reward is used to update the policy, iteratively improving the model’s reasoning capabilities across diverse domains.
In this work, we investigate reasoning domains beyond mathematics and analyze the impact of rl on llms trained with datasets from diverse domains and question formats. A core pre-requisite for effective self-learning is access to high-quality, diverse, and reward-compatible training data (Xie et al., 2025b; Hu et al., 2025). While mathematical reasoning has benefited from clean and verifiable datasets, extending rl to general-purpose reasoning domains remains underexplored due to the lack of structured, high-quality supervision. To address this, we explore methods for leveraging web documents and open-source QA benchmarks to collect general-purpose reasoning data. Incorporating a mix of structured and unstructured domains introduces a wide range of cognitive patterns and task-specific reasoning strategies which will further improve generalization. However, it introduces noise and ambiguity—particularly in open-ended formats—making it difficult to apply rule-based reward modeling reliably. To mitigate this, we apply task-specific templates to unify question and answer formats, limiting answer space variability and enabling simple but effective verifiable reward signals. Next, we apply a lightweight data filtering strategy to discard examples that are infeasible to verify—such as open-ended answers exceeding a certain length or mcqs with misaligned options—ensuring stable and interpretable rl training. Finally, we explore optimal data blending strategies that combine information across diverse domains and tasks. This allows us to investigate how the inclusion of general-purpose reasoning data complements mathematical reasoning, ultimately leading to broader and more adaptive generalization in llms.
Data Curation.
We start with carefully curating datasets from multiple sources to ensure diversity in the training data. Our training dataset $\mathcal{D}$ comprises two sources:
$\mathcal{D}=\mathcal{D}_{syn}\cup\mathcal{D}_{os}$
Here, $\mathcal{D}_{syn}\rightarrow$ synthetically generated data from Common Crawl (CC) and $\mathcal{D}_{os}\rightarrow$ publicly available open-source QA datasets. Each sources of data further consists of question answer pairs related to general purpose reasoning and mathematics:
$\mathcal{D}_{syn}\rightarrow\mathcal{D}_{syn\_gpr}\cup\mathcal{D}_{syn\_mr};\quad\mathcal{D}_{os}\rightarrow\mathcal{D}_{os\_gpr}\cup\mathcal{D}_{os\_mr}$
Data Source Category Type Samples
mmlu [Train] gpr mcq 99,842
Syn-qa gpr mcq 192,930
Natural Reasoning gpr oe 100,000
NuminaMath mr oe 87,350
PersonaSkill-math mr oe 100,000
Math mr oe 8523
Total 588,645
Table 1: Training data distribution by source and type. oe=Open-Ended; gpr =General-Purpose Reasoning; mr =Math Reasoning
  • General Purpose Reasoning, $\mathcal{D}_{gpr}$: We collect open source QA datasets ($\mathcal{D}_{os\_gpr}$)—Natural Reasoning (Yuan et al., 2025) and mmlu [Train] (Hendrycks et al., 2021a) that span multiple domains, including STEM fields (e.g., Physics, Computer Science), Economics, Social Sciences, and more. To enhance diversity, we further synthesize QA pairs from CC documents using the wide range of domains in mmlu as our seed domain. We denote this dataset as Syn-qa ($\mathcal{D}_{syn\_gpr}$).
    $\mathcal{D}_{gpr}\rightarrow\mathcal{D}_{syn\_gpr}\cup\mathcal{D}_{os\_gpr}$
  • Mathematical Reasoning, $\mathcal{D}_{mr}$: As mathematical questions inherently require Chain-of-Thought derivations which emphasizes the llm to think, we incorporate math reasoning corpus to our training data. We combine open-source mathematical reasoning datasets ($\mathcal{D}_{os\_mr}$), such as MATH (Hendrycks et al., 2021b) and Numina-Math (Beeching et al., 2024). We generate additional math problems applying the similar technique as Ge et al. (2024) and define it as Persona-math ($\mathcal{D}_{syn\_mr}$).
    $\mathcal{D}_{mr}\rightarrow\mathcal{D}_{syn\_mr}\cup\mathcal{D}_{os\_mr}$
Applying Templates for Answer Space and Reasoning Diversity.
General purpose reasoning benchmarks are often divided into two categories: (a) Multiple Choice Questions (Hendrycks et al., 2021a; Wang et al., 2024) and (b) Open-Ended Questions (Zhong et al., 2023). Recent works have ignored these variations in the answer space for consistent reward design across all tasks which are often predominantly math tasks (Hu et al., 2025; Aggarwal & Welleck, 2025; Luo et al., 2025). We hypothesize that each question type elicits different thinking patterns, leading to diverse reasoning trajectories in the model. Training on different question types will enhance the model’s ability to generalize by exposing it to diverse answer formats, thereby fostering different reasoning pathways.
Therefore, to observe the effect of question type in rl training, we synthesize $\mathcal{D}_{gpr}$ using two templates: $\mathcal{T}_{MCQ}$ Multiple Choice Questions (mcq), and $\mathcal{T}_{Open}$ Open-Ended questions. We convert the mcq datasets (mmlu) to open-ended by removing the options from the questions.
$\mathcal{D}_{mcq}=\mathcal{T}_{MCQ}(\mathcal{D}_{gpr}),\quad\mathcal{D}_{open}=\mathcal{T}_{Open}(\mathcal{D}_{gpr})$
Additionally, some mcq questions are incomplete without options (e.g., Which of the following ways we can file taxes?). We discard such questions to avoid confusion during answer generation. Finally, our general purpose reasoning data, $\mathcal{D}_{gpr}$, can be represented as:
$\mathcal{D}_{gpr}=\mathcal{D}_{mcq}\cup\mathcal{D}_{open}$
Data Filtering and Formatting.
To ensure high-quality training data, we apply a series of filtering and formatting steps, $\mathcal{H}$, to remove samples that are infeasible to evaluate with a simple rule-based reward function. Specifically, for $\mathcal{D}_{mcq}$, we check whether the correct answer appears within the question text itself. Given a question-answer pair $(q,a^{*})$ with answer choices $\{a_{1},a_{2},\dots,a_{n}\}$, we discard a sample if $a^{*}\notin\{a_{1},a_{2},\dots,a_{n}\}$.
For $\mathcal{D}_{open}$, such as those in the Natural Reasoning dataset, we discard samples that are challenging to evaluate with a rule-based reward function. Formally, we retain samples where $|w(a^{*})|\leq 10$; $w(a^{*})$ represents the number of words in the answer $a^{*}$.
Lastly, for the mathematical reasoning corpus, $\mathcal{D}_{mr}$, we remove entries that lack an associated answer, ensuring that all retained questions $q$ have a valid response $a^{*}$, i.e., we discard samples where $a^{*}=\emptyset$.
$\mathcal{D}^{\prime}=\mathcal{H}(\mathcal{D})=\left\{(q,a^{*},\{a_{1},\dots,a_{n}\})\in\mathcal{D}\;\middle|\;\begin{array}[]{l}a^{*}\in\{a_{1},\dots,a_{n}\}\quad(\mathcal{D}_{mcq})\\ |w(a^{*})|\leq 10\quad(\mathcal{D}_{open})\\ a^{*}\neq\emptyset\quad(\mathcal{D}_{mr})\end{array}\right.$
Data Blending.
We study the impact of data diversity in three paradigms:
Category Blend Name Symbol Blend Description
Data Source Natural Distribution $\mathcal{B}_{nd}$ Ratio of number of samples in a dataset divided
by the total number of samples in all the datasets.
More Math $\mathcal{B}_{mr\uparrow}$ 2:1 ratio of $\mathcal{D}_{mr}$ and $\mathcal{D}_{gpr}$
More General Purpose Reasoning $\mathcal{B}_{gpr\uparrow}$ 2:1 ratio of $\mathcal{D}_{gpr}$ and $\mathcal{D}_{mr}$
Question Types More mcq $\mathcal{B}_{mcq\uparrow}$ 2:1 ratio of $\mathcal{D}_{mcq}$ and $\mathcal{D}_{open}$
More Open-Ended $\mathcal{B}_{open\uparrow}$ 2:1 ratio of $\mathcal{D}_{open}$ and $\mathcal{D}_{mcq}$
Data Usefulness Avg. Score $\mathcal{B}_{score}$ Provide weight to each source based
on their average benchmark performances
Table 2: Overview of Data Blending Strategies. Blends are categorized by data source, question type, and usefulness—each constructed to assess the impact of domain diversity, format variation, and task relevance on RL-based reasoning.
  • Data Source: We have gathered questions from diverse domains including math ($\mathcal{D}_{mr}$), STEM, humanities, economics, history, law, social sciences, etc., ($\mathcal{D}_{gpr}$) and observe the effect of each source on rl training.
  • Question Types: We investigate the impact of question types in downstream tasks.
  • Data Usefulness: We further analyze what is the contribution of each data sources in downstream task performances. We initially run RL using individual data alone and then evaluate them across diverse downstream tasks. Based on their performances, we create a new blend.
Based on these three categories, we construct six distinct blends, summarized in Table 2, with their corresponding dataset weight distributions detailed in Table 8.
Reinforcement Learning with grpo.
We begin with a pretrained large language model (llm) $\mathcal{M}$ and a training blend $\mathcal{B}$, where each sample contains only the input prompt and the final answer which is verifiable. We employ Group Relative Policy Optimization (grpo) (Shao et al., 2024). grpo does not use a separate critic model and instead estimates the baseline from group scores, improving efficiency and reducing memory. For each question $q$, grpo samples a group of outputs $o_{1},o_{2},...,o_{G}$ from the old policy $\pi_{\theta_{old}}$ and then optimizes the policy model $\pi_{\theta}$ by maximizing the following objective:
$\mathcal{J}_{\text{{grpo}}}(\theta)=$ $\;\mathbb{E}\left[q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O|q)\right]$
×$\times\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\Bigg{[}\min\Bigg{(}\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}\hat{A}_{i,t},\text{clip}\Bigg{(}\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})},1-\epsilon,1+\epsilon\Bigg{)}\hat{A}_{i,t}\Bigg{)}$
$\quad-\beta D_{\text{KL}}\big{(}\pi_{\theta}\|\pi_{\text{ref}}\big{)}\Bigg{]}$
$D_{\text{KL}}\left[\pi_{\theta}\|\pi_{\text{ref}}\right]=\frac{\pi_{\text{ref}}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}-\log\frac{\pi_{\text{ref}}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}-1.$ (1)
where $\epsilon$ and $\beta$ are hyperparameters, and $\hat{A}_{i,t}$ is the advantage, computed using a group of rewards $\{r_{1},r_{2},...,r_{G}\}$ corresponding to the outputs within each group:
$\hat{A}_{i,t}=\frac{r_{i}-\text{mean}(\{r_{1},r_{2},...,r_{G}\})}{\text{std}(\{r_{1},r_{2},...,r_{G}\})}$
Rule Based Reward Modeling.
To guide the reinforcement learning process, we employ a rule-based reward system designed for verifiable evaluation. Similar to (DeepSeek-AI, 2025), we define the total reward function $\mathcal{R}$ as the logical and of an accuracy reward $\mathcal{R}_{\text{acc}}$ and a format reward $\mathcal{R}_{\text{format}}$:
$\mathcal{R}=\mathcal{R}_{\text{acc}}\wedge\mathcal{R}_{\text{format}}.$
This implies that the output will get reward only when both the answer and the format are correct.
Accuracy Reward: The accuracy reward evaluates correctness based on whether the model’s response $p$ is similar to the ground truth solution $a$ to satisfy the correctness criteria:
$\mathcal{R}_{\text{acc}}(p,a)=\begin{cases}1,&\text{if equal}(p,a),\\ 0,&\text{otherwise}.\end{cases}$
[…]
$R_{\text{format}}(a)=\begin{cases}1,&\text{if }F(a),\\ 0,&\text{otherwise}.\end{cases}$
where $F(a)$ returns True if $a$ is correctly formatted and False otherwise.
3 Experimental Setup
Training Details.
We adopt Qwen2.5-7B and Qwen2.5-32B (Team, 2024a) as our baseline models, $\mathcal{M}$, which demonstrate strong generalization capabilities across various natural language reasoning tasks. We directly apply grpo training on $\mathcal{M}$ using the veRL framework [^1] https://github.com/volcengine/verl, which is an open-source implementation of the HybridFlow RLHF framework (Sheng et al., 2024). We train the base models with key settings including a constant learning rate of 1e-6, a batch size and PPO mini batch size of 128 and a maximum context length of 5000 tokens. Each generation step contains 128 unique prompts sampled from the dataset, and performing 8 rollouts with temperature and top-p both set to 1.0. We set KL coefficient to 0.001 in all experiments. During training, the model is directly exposed to mixed types of questions from different domains. Note that we did not conduct extensive hyperparameter tuning, so one can expect further improvements with additional optimization.
Evaluation Metrics.
[…]
4 Experiments and Results
Analyze the effect of Individual Datasets.
To prepare an effective blend using diverse sources of data, we begin by understanding impact of individual data sources on the self-learning paradigm so that we can prioritize the useful data sources and provide less weights to others. In this setup, we employ self-learning using $\mathcal{M}$=Qwen-2.5-7B and taking each dataset separately. To make consistent comparison across different data sources, we keep the training recipe constant for all experiments. We run a controlled experiments and train each models for fewer steps (250 steps) and evaluate them on the last checkpoint.
Data Source mmlu mmlu-pro gpqa-diamond agieval supergpqa math-500 amc23 Avg
$\mathcal{M}$ 74.20 45.00 31.82 48.59 25.36 48.30 40.00 44.75
mmlu [Train] 69.76 38.50 32.83 47.66 27.69 22.00 5.00 34.78
Syn-qa 70.45 52.41 30.81 52.10 24.57 54.20 35.00 45.65
Natural Reasoning 68.89 31.33 33.33 46.65 22.44 68.60 42.50 44.82
NuminaMath 72.94 52.05 33.84 54.39 26.97 76.20 55.00 53.06
PersonaSkill-Math 53.99 28.08 18.69 45.69 16.92 77.20 50.00 41.51
Math 63.30 31.64 21.72 51.95 18.31 78.40 50.00 45.04
Table 3: Results of Self-Learning on Individual Datasets. Each row shows the downstream evaluation results after self-learning on a single data source. Results highlight the varying strengths of individual datasets across general-purpose and mathematical benchmarks.
Table 2 shows that different datasets have varying impacts on downstream accuracies across reasoning benchmarks. Notably, the NuminaMath yields the highest overall average, outperforming the baseline ($\mathcal{M}$) by over 8.30%. Its strength is especially pronounced on mathematical tasks such as math-500 and amc23 but additionally it achieves superior accuracies on general purpose reasoning tasks showing a strong generalization across diverse domains. The Syn-qa dataset demonstrates a $\sim$1.0improvement over baseline with stronger accuracy in mmlu-pro, agieval and math-500 tasks, suggesting that synthetically generated instruction-style data can generalize well when aligned with benchmark distributions. Natural Reasoning, despite modest scores on language-rich benchmarks, delivers a surprisingly strong overall average, driven by high scores in math-500 and amc23. This indicates that reasoning-focused datasets, even if less curated, can contribute meaningfully in math-adjacent tasks. On the other hand, Persona-Math, although strong in math, suffers from low generalization across most benchmarks. Finally, the mmlu [Train] dataset underperforms across most tasks, specifically in math reasoning domains, suggesting that self-learning with raw mmlu [Train] data alone is insufficient for generalization. However, it obtains the best score for supergpqa, which requires reasoning across wide range of cross-disciplinary domains. This highlights the potential of mmlu [Train] in capturing broad conceptual knowledge and supporting transfer to long-tail domains, making it a valuable component when targeting general-purpose reasoning benchmarks. While preparing blends for Data Usefulness, we use the average accuracies of individual sources to obtain $\mathcal{B}_{score}$ i.e., we provide more weight to datasets like Syn-qa, NuminaMath and less to mmlu [Train].
Analysis across Blends.
We observe the effect of Nemotron-CrossThink in three different categories using six different blends. To show the distinction between natural distribution and selective weighting of domains, we also prepare $\mathcal{B}_{nd}$, which represents data sampled in proportion to each dataset’s original size. Additionally, to analyze the impact of within-domain training versus cross-domain blending, we introduce a separate category called Single Source. We prepare two domain-specific blends: $\mathcal{B}_{only\_mr}$, using only $\mathcal{D}_{mr}$ data, and $\mathcal{B}_{only\_gpr}$, using only $\mathcal{D}_{gpr}$ data. We further compare Nemotron-CrossThink with a recent math-centric self-learning approach, Open-Reasoner-Zero (orz) (Hu et al., 2025)—which achieved superior accuracy in math benchmarks by training rl on combination of math data. For fair comparison we evaluate the 7B model using our eval setup.
Model Category Blend mmlu mmlu-pro gpqa-diamond agieval supergpqa math-500 amc23 Avg
$\mathcal{M}$ 74.20 45.00 31.82 48.59 25.36 48.30 40.00 44.75
orz 73.20 48.90 29.30 63.49 27.60 81.40 62.50 55.20
*CrossThink $\mathcal{B}_{nd}$ 73.18 54.81 38.07 59.99 26.54 77.00 60.00 55.66
Data Source $\mathcal{B}_{mr\uparrow}$ 74.85 55.51 40.10 61.47 26.81 77.80 67.50 57.72
$\mathcal{B}_{gpr\uparrow}$ 74.94 57.82 38.58 63.71 29.16 77.60 65.00 58.12
Question Types $\mathcal{B}_{mcq\uparrow}$ 74.26 55.77 39.59 62.54 28.05 78.00 60.00 56.89
$\mathcal{B}_{open\uparrow}$ 74.46 55.82 43.15 61.28 26.82 78.40 62.50 57.49
Data Usefulness $\mathcal{B}_{score}$ 74.70 56.16 40.10 59.80 27.37 78.00 62.50 56.95
Single Source $\mathcal{B}_{only\_mr}$ 74.24 54.26 38.58 61.39 27.69 78.60 70.00 57.82
$\mathcal{B}_{only\_gpr}$ 72.77 52.06 37.06 56.56 27.44 72.20 55.00 53.30
Table 4: Results of Nemotron-CrossThink-7B across Blends. Multi-domain blend $\mathcal{B}{gpr\uparrow}$ achieves the highest overall average accuracy, outperforming domain-specific and naturally sampled blends—underscoring the benefit of self-learning with diverse reasoning data. (*) Due to the space shortage, we use *CrossThink to refer Nemotron-CrossThink.
As shown in Table 4, each blending strategy consistently outperforms the base model, $\mathcal{M}$, by a significant margin. The natural distribution blend, $\mathcal{B}_{nd}$, yields a notable improvement of over 13% on average compared to $\mathcal{M}$, suggesting that simply increasing the amount of training data—even without rebalancing—can be beneficial.
$\mathcal{B}_{gpr\uparrow}$ from the Data Source category achieves the highest overall average, as well as the strongest results across most reasoning-focused benchmarks (e.g., +12.82% on mmlu-pro and +15.12% on agieval). Notably, it performs relatively $\sim 5\%$ on average better than orz. While $\mathcal{B}_{only\_math}$ performs slightly better on math-specific tasks, such as a marginal 1% gain on math-500, it lags behind on non-math reasoning benchmarks—underperforming $\mathcal{B}_{gpr\uparrow}$ by $\sim$3–4on tasks like agieval, supergpqa, and mmlu-pro. The same trend is also seen with orz. To better understand these differences, we analyze sub-category accuracies in Appendix C and find that $\mathcal{B}_{gpr\uparrow}$ shows large relative gains in non-math categories while differences in math subcategories are either negligible or even favor $\mathcal{B}{gpr\uparrow}$ in some tasks. This highlights that general-purpose reasoning data offers strong cross-domain transfer with minimal compromise on math accuracy, making it more versatile.
Both $\mathcal{B}_{mcq\uparrow}$ and $\mathcal{B}_{open\uparrow}$ in Question Types category show consistent gains, with the latter achieving a slight edge (0.6% improvement on average). In addition, $\mathcal{B}_{open\uparrow}$ yields stronger results on mathematical benchmarks. Mathematical problems are inherently open-ended in structure. As a result, highlighting more open-ended domains aligns with the format and reasoning demands of math tasks. This suggests that diversity in question formats—especially open-ended reasoning—can better generalize to both general purpose reasoning and math-focused downstream tasks.
Regarding Data Usefulness, the score-based selection strategy ($\mathcal{B}_{score}$) outperforms the base model $\mathcal{M}$, indicating the effectiveness of selective data curation. However, despite focusing more on the better performing datasets in Table 3, $\mathcal{B}_{score}$ is overall worse than blends like $\mathcal{B}_{mr\uparrow}$ or $\mathcal{B}_{only\_math}$. This gap arises because $\mathcal{B}{score}$ assigns weights based solely on average dataset scores, without accounting for task-specific strengths. For instance, Math and Persona-Math receive higher weights than Natural Reasoning or MMLU due to their math accuracy, despite the latter performing significantly better on general-purpose reasoning tasks. In contrast, domain-aware blends selectively prioritize datasets based on their utility within specific domains, leading to more effective coverage and stronger scores across both math and general-purpose reasoning tasks.
To investigate the impact of single-domain versus mixed-domain training data in rl, we compare the Single Source category with other blending strategies. Notably, $\mathcal{B}_{only\_mr}$ achieves the highest average math score (56.20%) among all blends, ranking as the second-best blend overall in terms of average accuracy. In contrast, while $\mathcal{B}_{only\_gpr}$ outperforms the base model $\mathcal{M}$, it underperforms in mathematical reasoning tasks. Surprisingly, despite being tailored for general-purpose reasoning, $\mathcal{B}_{only\_gpr}$ also lags behind $\mathcal{B}_{only\_mr}$ by 4.2% on average across non-math reasoning benchmarks. This counterintuitive finding suggests that to obtain maximum gain in general purpose reasoning tasks we need to include mathematical problems in the training blend. As discussed earlier, $\mathcal{B}_{gpr\uparrow}$ gets the best average reasoning accuracy which consists of both math and general purpose reasoning datasets. This confirms that math data alone is transferable to structured reasoning tasks, whereas general-purpose data is less effective when isolated.
5 Ablations
Nemotron-CrossThink is token efficient in responses
To further understand the influence of multi-domain data in response generation, we compare the average token lengths of correct and incorrect responses between models trained on two blends: $\mathcal{B}_{gpr\uparrow}$ and $\mathcal{B}_{only\_mr}$. As shown in Figure 3, on general-purpose reasoning (gpr) benchmarks, $\mathcal{B}_{gpr\uparrow}$ consistently outperforms $\mathcal{B}_{only\_mr}$ and orz (Hu et al., 2025), not only in accuracy (as shown in Table 4) but also in response efficiency—producing correct answers with significantly fewer tokens [^2] Detailed categorization per task is shown in Appendix B.. For instance, on mmlu, the average token count for correct responses is 229 for $\mathcal{B}_{gpr\uparrow}$, compared to 351 for $\mathcal{B}_{only\_mr}$. This demonstrates that exposure to multi-domain data enables the model to internalize a more efficient reasoning strategy, leading to both improved performance and reduced inference cost.
Figure 3: Token efficiency comparison of models trained on $\mathcal{B}_{gpr\uparrow}$ (multi-domain blend) and two single domain blends ($\mathcal{B}_{only\_math}$ and orz).
In contrast, on math-specific benchmarks, $\mathcal{B}_{only\_mr}$ and orz perform slightly better in accuracy, as expected due to domain alignment. Interestingly, correct responses are generally longer than in reasoning tasks as solving math problems inherently requires detailed, multi-step derivations, hypothesis exploration, verification and refinement. Despite this, the $\mathcal{B}_{gpr\uparrow}$ shows its adaptability by generating longer responses for math tasks and shorter ones for gpr tasks—indicating a dynamic response strategy learned through multi-domain training. As shown in Table 9, $\mathcal{B}_{gpr\uparrow}$ has a wide dynamic range for generating responses. It increases its average tokens by 62% when generating responses for math tasks (Mean Tokens=622) as opposed to general reasoning tasks (Mean Tokens=385). Whereas, $\mathcal{B}_{only\_mr}$ increases it average tokens only by 14% (Mean Tokens=731 for math tasks and Mean Tokens=639 for general reasoning tasks) showing a much smaller dynamic range. This trend is also mirrored in orz, trained on a high-quality blend of math datasets, which shows an even smaller increase (12%) in average token length across domains.
This adaptive behavior highlights a key strength of multi-domain training: it equips the model with the flexibility to tailor its response style to the nature of the task. By learning from a diverse range of domains, $\mathcal{B}_{gpr\uparrow}$ learns to reason efficiently—across all tasks, $\mathcal{B}_{gpr\uparrow}$ uses on average 28% fewer tokens for correct responses than $\mathcal{B}_{only\_mr}$—producing compact yet accurate answers where appropriate, and detailed ones when necessary.
Data Format Study: Question and Answer Templates.
To better understand how training data formatting affects model performance, we conduct two controlled studies focused on question and answer template design, as shown in Table 5 and Table 6.
In Table 4, we observe that $\mathcal{B}_{open\uparrow}$ outperforms $\mathcal{B}_{mcq\uparrow}$, suggesting that models trained on more open-ended data generalize better across benchmarks. This motivated us to investigate whether converting all questions into a unified open-ended format leads to better performance. In Question Template Study, we use the natural distribution blend ($\mathcal{B}_{nd}$) and only perturb the question template . To generate the open-ended variant, we remove the answer options from mcqs, prompting the model to produce an answer without selecting from predefined choices.
Question Type mmlu mmlu-pro gpqa-diamond agieval supergpqa math-500 amc23 Avg
mcq +open-ended 73.18 54.81 38.07 59.99 26.54 77.00 60.00 55.66
open-ended 74.61 54.36 39.09 59.30 29.16 76.60 65.00 56.87
Table 5: Impact of Question Format. Converting all questions to open-ended format improves accuracy across benchmarks, reducing reliance on option guessing and encouraging deeper reasoning.
Table 5 illustrates that the open-ended-only configuration consistently outperforms the mixed-format setting across nearly all benchmarks, achieving 1.21% higher average score. Notably, it leads to significant improvements on reasoning-intensive and traditionally mcq-formatted benchmarks such as mmlu, supergpqa, and gpqa-diamond. This result may be attributed to the inherent structure of mcq questions, where random guessing can yield an accuracy of approximately 25% in mmlu and gpqa-diamond benchmarks where we have only four options. In contrast, open-ended questions eliminate this guessing advantage, compelling the model to rely more heavily on reasoning to arrive at a correct answer. By reducing the likelihood of reward hacking through random option selection, the open-ended format encourages more robust reasoning and leads to improved generalization.
In the Answer Template Study, we investigate how the format of output labels influences training effectiveness on mcq-style datasets. We compare two answer templates: Long - the model is trained to generate both the option label and its corresponding description (e.g., (A) The sky is blue), and Short - the model is trained to output only the option label (e.g., A). For this study, we use the $\mathcal{B}_{only\_gpr}$ blend, which primarily consists of mcq datasets (Table 1), making it ideal for analyzing the effects of answer formatting in this setting.
Answer Type mmlu mmlu-pro gpqa-diamond agieval supergpqa math-500 amc23 Avg
Long 72.77 52.06 37.06 56.56 27.44 72.20 55.00 53.30
Short 74.22 54.56 39.59 58.01 28.39 74.20 52.50 54.50
Table 6: Impact of Answer Format. Using short-form answers improves accuracy by reducing output ambiguity and avoiding penalization from rigid reward functions in rule-based training.
Table Table 6 shows that the short-form answer template consistently outperforms the long-form variant, with a 1.20% improvement in average accuracy. This trend holds across both reasoning and mathematical benchmarks. These results suggest that reducing the complexity of the output space helps minimize ambiguity and allows the model to better align its predictions with the structure of the question. Furthermore, when training with long-form answers using a rule-based reward (e.g., exact string matching), the model is frequently penalized for minor deviations in phrasing, even when the correct option is selected. For instance, if the model outputs the correct option label but paraphrases the description slightly, the strict reward signal treats it as incorrect. This introduces noisy supervision and may hinder learning. While this issue could be mitigated by designing a more flexible reward function (e.g., based on semantic similarity or option-label matching), our goal in this work is to keep the approach simple and interpretable. As such, we adopt a naive rule-based reward for clarity and reproducibility, and leave more sophisticated reward designs for future investigation.
Difficulty Filtering.
Training with high-quality data is a key factor in self-learning to ensure efficient and stable learning and to obtain correct reward signals. Recent works (Hu et al., 2025; Luo et al., 2025; Cui et al., 2025) have explored various filtering strategies to remove noisy reference answers from datasets, focusing on data that is easily verifiable using simple rule-based rewards. Zeng et al. (2025) further investigate data selection based on question complexity, showing that as the difficulty of the training data increases, the resulting model achieves better downstream accuracy. However, their approach relies on datasets like math-500 that come with predefined difficulty scores. In this work, we explore a simple approach to estimate question difficulty for general purpose reasoning datasets that do not come with explicit difficulty labels. Specifically, we label questions as ‘difficult’ if they are answered incorrectly by a smaller model (Qwen-2.5-7B) in a zero-shot setting and filter out the ‘easy’ questions. The intuition is that questions easily answered by a base model are likely to be knowledge-based or shallow in reasoning depth, whereas those it fails on are likely to require deeper reasoning or broader generalization. We construct two versions of our training dataset $\mathcal{B}_{gpr\uparrow}$—an unfiltered set containing all questions, and a filtered set ($\mathcal{B}_{f(gpr)\uparrow}$) that retains only the difficult samples—and use them to train separate instances of a larger model $\mathcal{M}$ = Qwen-2.5-32B.
Model Blend mmlu mmlu-pro gpqa-diamond agieval supergpqa math-500 amc23 Avg
Qwen-2.5-32B 83.30 55.10 40.40 62.77 33.16 60.55 45.00 54.33
Nemotron-CrossThink-32B $\mathcal{B}_{gpr\uparrow}$ 83.57 68.83 46.70 73.90 37.99 82.40 67.50 65.84
$\mathcal{B}_{f(gpr)\uparrow}$ 83.60 69.43 49.75 75.82 38.34 84.00 75.00 67.99
Table 7: Difficulty-Based Filtering. Filtering $\mathcal{B}_{gpr\uparrow}$ to retain only hard examples ($\mathcal{B}_{f(gpr)\uparrow}$) yields consistent gains across all tasks, highlighting the effectiveness of selective training on challenging data.
According to Table 7, this filtering approach results in consistent performance improvements across all evaluated benchmarks. While both filtered and unfiltered models outperform the original baseline Qwen-2.5-32B, the model trained on the filtered dataset—denoted as $\mathcal{B}_{gpr\uparrow}{\text{f}}$—achieves the highest accuracy on every task. The gains are especially prominent in complex benchmarks such as mmlu-pro, gpqa-diamond, agieval, and amc23, where the filtered model improves by up to 2–8% over its unfiltered counterpart. On average, filtering boosts overall accuracy by 2.15%, a notable gain considering that it comes from training on fewer but harder examples. This suggests that selectively training on challenging examples can yield more robust and generalizable models, likely due to stronger gradient signals and a focus on harder-to-learn reasoning patterns.
6 Related Work
Evolution of Reasoning in llm.
Large Language Models have demonstrated remarkable dominance across numerous Natural Language Processing tasks. To enhance the complex reasoning capabilities of llms, (Wei et al., 2022) introduce Chain-of-Thought (CoT), which incorporates multi-step intermediate reasoning before arriving at final conclusions. CoT exhibits significant advantages across multiple domains, including mathematics, science, and programming. Subsequently, (OpenAI, 2024) further explore CoT and propose the Long Chainof-Thought framework. In Long CoT, llms demonstrate advanced cognitive behaviors such as reflection, verification, correction, and multipath exploration, thereby further enhancing their problem-solving capabilities in complex reasoning tasks. Moreover, Long CoT exhibits excellent test-time scaling properties, where increased computational resources correlate with improved reasoning outcomes. Models like QwQ (Team, 2024b; 2025b), DeepSeek-R1 (DeepSeek-AI, 2025), Kimi k1.5 (Team, 2025a), and InternThinker (Cai et al., 2024) have successfully experimented with Long CoT for enhanced reasoning, combining fine-tuning and Reinforcement Learning to elevate the performance of open-source reasoning models to unprecedented levels. Notably, subsequent models such as Open-Reasoner-Zero (Hu et al., 2025), Open-R1 (Face, 2025), O1-Replication (Qin et al., 2024; Huang et al., 2024; 2025), s1 (Muennighoff et al., 2025) and LIMO (Ye et al., 2025) observes significant benefits from Long CoT even in smaller models through simple distillation.
Self-Learning beyond Math.
High-quality training data are crucial for scalable Reasoner-Zero training. Most of the recent works emphasize mathematical benchmark-centric data (AMC, AIME, Math, Olympiads, and AoPS) for reinforcement learning (Hu et al., 2025; Aggarwal & Welleck, 2025; Trung et al., 2024; Ye et al., 2025; Zeng et al., 2025) as designing verifiable rewards is much easier for math tasks. They exclude problems such as multiple choice and proof-oriented problems which reduces the answer space diversity. mcq type of questions are important for mmlu and other non-reasoning centric tasks. For a rule-based reward model, the format of input data and the final answer is crucial and largely underexplored. Furthermore, their additional sources of data synthesis approach has no details making it infeasible to scale for domains other than math. The kind of data and the ratio of each type of data important for the overall improvement of llms across multiple benchmarks have yet to be explored.
Data Sampling in rl.
Recent works have widely explored the idea of combining data from multiple sources during rl training to enhance the diversity of reasoning tasks and improve model generalization (Hu et al., 2025; Luo et al., 2025; Zeng et al., 2025; Wen et al., 2025). These studies primarily concentrate on the mathematical domain, where rule-based correctness allows for straightforward reward modeling. In such setups, data sampling strategies are often driven by factors like question complexity or the ease with which answers can be verified algorithmically. For instance, questions are filtered or prioritized based on whether they are solvable with deterministic programs or satisfy certain symbolic constraints. A notable direction is curriculum learning, where Xie et al. (2025b) utilizes synthetically generated puzzle-like data from Xie et al. (2025a) to control the difficulty level and study the progression of learning. However, these works remain narrowly focused on highly structured domains such as logic puzzles or math word problems. Yeo et al. (2025) has shown that including 50% of math and 50% of noisy verifiable data from WebInstruct-462k Yue et al. (2024) yields best mmlu-pro score in rl setup—indicating the potential of mixing of domains in the training blend. However, it in unclear how this benefit is attributed to the inclusion non-math reasoning data as 68.36% of WebInstruct-462k is about math. They have performed filtering to obtain data with feasible verifiable reward and this is boost and prioritize the mathematical domain over other domains. Despite this progress, there is a lack of systematic investigation into how including non-math reasoning data—such as legal analysis, social science, commonsense inference, or historical interpretation—affects rl training. Nemotron-CrossThink is the first systematic framework to incorporate multi-domain and multi-format data into rl, introducing verifiable reward mechanisms for non-deterministic domains and demonstrating that blending diverse reasoning sources leads to stronger generalization across benchmarks.
7 Conclusion
We present Nemotron-CrossThink, a simple and scalable framework for improving the generalization abilities of LLMs through reinforcement learning with multi-domain corpora. By combining data from diverse reasoning domains and applying lightweight filtering and formatting strategies, Nemotron-CrossThink enables consistent gains across both general-purpose and mathematical benchmarks. Our best-performing blend—constructed with a 2:1 ratio of general-purpose to math data—achieves a 13.36% average improvement over strong baselines, with gains further amplified by difficulty-based filtering and thoughtful template design. Importantly, these benefits persist across model scales and task types, demonstrating that data diversity, not just data volume, is key to broader reasoning capabilities. Nemotron-CrossThink offers a practical recipe for building more generalizable, efficient, and reliable LLMs under the rl paradigm—paving the way for scalable self-learning beyond math.
Appendix A Data Proportion across Blends
To better understand the data composition used in our reinforcement learning experiments, we report the proportion of each dataset in the six blending strategies introduced in Section 2. These proportions reflect how data is distributed across different sources depending on the specific blending paradigm: data source, question type, and data usefulness.
Data Name Type $\mathcal{B}_{nd}$ $\mathcal{B}_{mr\uparrow}$ $\mathcal{B}_{mcq\uparrow}$ $\mathcal{B}_{open\uparrow}$ $\mathcal{B}_{gpr\uparrow}$ $\mathcal{B}_{score}$ $\mathcal{B}_{only\_math}$ $\mathcal{B}_{only\_gpr}$
mmlu MCQ 0.1696 0.0864 0.2251 0.1159 0.1678 0.1296 0.2542
Syn-qa MCQ 0.3277 0.1670 0.4349 0.2241 0.3242 0.1731 0.4912
Natural Reasoning open-ended 0.1699 0.0866 0.1149 0.2231 0.1680 0.1683 0.2546
NuminaMath open-ended 0.1484 0.2943 0.1004 0.1949 0.1516 0.2020 0.4460
Persona-math open-ended 0.1699 0.3370 0.1149 0.2231 0.1736 0.1579 0.5105
math open-ended 0.0145 0.0287 0.0098 0.0190 0.0148 0.1691 0.0435
Table 8: Proportion of each dataset in different blends.
Appendix B Token Efficiency Analysis
Token Efficiency in Correct Responses.
Understanding not only whether a model answers correctly but also how efficiently it reasons is critical in real-world deployments, especially for reducing inference cost and latency. To this end, we analyze the token lengths of correct responses generated by models trained under different data blending strategies.
Table 9 presents the minimum, maximum, and mean number of tokens used in correct answers across two task types: General Purpose Reasoning (GPR) and Math. We compare three models: (1) $\mathcal{B}_{gpr\uparrow}$ (multi-domain training), (2) $\mathcal{B}_{only\_math}$ (math-only training), and (3) orz (a strong math-centric baseline model).
Task Type Model Min Max Mean
GPR $\mathcal{B}_{gpr\uparrow}$ 83.20 2697.80 385.41
$\mathcal{B}_{only\_math}$ 159.60 9594.00 638.57
orz 223.00 8221.80 1114.60
Math $\mathcal{B}_{gpr\uparrow}$ 170.25 10130.00 622.00
$\mathcal{B}_{only\_math}$ 201.75 11330.25 730.68
orz 292.00 12917.00 1257.00
Table 9: Token length statistics (Min, Max, Mean) for correct responses across task types.
Across gpr tasks, $\mathcal{B}_{gpr\uparrow}$ produces the most concise correct responses, with a mean of 385 tokens—39.6% fewer than $\mathcal{B}_{only\_math}$ and 65.4% fewer than orz. This suggests that training with multi-domain corpora equips the model to reason more efficiently in less structured tasks, avoiding unnecessarily verbose responses.
On math benchmarks, where detailed step-by-step derivations are essential, all models naturally generate longer outputs. However, $\mathcal{B}_{gpr\uparrow}$ still demonstrates adaptability, producing appropriately longer responses compared to GPR, while keeping the output concise relative to $\mathcal{B}_{only\_math}$ and orz. This behavior underscores the ability of multi-domain trained models to dynamically adjust their reasoning strategy and verbosity based on task requirements.
Interestingly, orz exhibits the longest response lengths across both GPR and math tasks. While this aligns with its design as a reasoning-heavy model, it also reflects less efficiency—potentially generating unnecessarily long chains of thought, particularly in domains outside its training focus.
In summary, the token efficiency analysis reveals that $\mathcal{B}_{gpr\uparrow}$ achieves a favorable trade-off between accuracy and brevity, tailoring its reasoning depth to the complexity of the task. This reinforces the value of diverse, multi-domain training in promoting adaptable and cost-efficient language models.
Figure 4: Average token lengths of correct and incorrect responses across general-purpose and math reasoning tasks for models trained on $\mathcal{B}_{gpr\uparrow}$, $\mathcal{B}_{only\_math}$, and orz.
Thinking Long vs Thinking Accurate.
Recent studies such as DeepScaler (Luo et al., 2025) have noted that incorrect answers often exhibit longer trajectories, leading to wasted computation and less efficient learning. Echoing this observation, we analyze the average token lengths of correct and incorrect responses for models trained on different blends: $\mathcal{B}_{gpr\uparrow}$, $\mathcal{B}_{only\_math}$, and orz.
As shown in Figure 4, incorrect responses are consistently and substantially longer than correct ones—by 3.6$\times$ on average. This pattern holds across both general-purpose and math reasoning tasks, suggesting that verbose reasoning does not guarantee correctness. In fact, longer responses often reflect the model’s uncertainty, overthinking, or repetitive CoT traces, rather than productive deduction.
Appendix C Sub-category Accuracy Analysis
To further support our observation that multi-domain training improves general-purpose reasoning while remaining competitive on math tasks, we analyze the number of correct responses across sub-categories in mmlu-pro and agieval. Figure 5 and Figure 6 show the count of correct answers produced by $\mathcal{B}_{gpr\uparrow}$ and $\mathcal{B}_{only\_math}$ across their respective sub-domains.
Figure 5: Sub-category Accuracy Comparison across mmlu-pro Domains. The $\mathcal{B}_{gpr\uparrow}$ blend consistently outperforms $\mathcal{B}_{only\_math}$ in a wide range of non-math reasoning categories such as business, law, psychology, and economics. Surprisingly, it also slightly surpasses the math-specialized blend in the mmlu-pro math category, highlighting the generalizability and versatility of multi-domain training.
Figure 6: Sub-category Accuracy Comparison across agieval. While $\mathcal{B}_{only\_math}$ performs marginally better in the math, $\mathcal{B}_{gpr\uparrow}$ achieves stronger results in non-math domains.
On mmlu-pro, $\mathcal{B}_{gpr\uparrow}$ consistently outperforms $\mathcal{B}_{only\_math}$ across non-math reasoning categories such as business, law, psychology, chemistry, and economics. Notably, it achieves relative improvements of +20.58% in law and +13.26% in business. Surprisingly, $\mathcal{B}_{gpr\uparrow}$ also performs better in the math category (+7.2%), despite not being trained exclusively on mathematical data. This may be attributed to the nature of mmlu-pro’s math problems, which are college-level and benefit from a combination of symbolic and heuristic reasoning—skills reinforced through exposure to diverse domains.
In contrast, the agieval benchmark (shown in Figure 6) features Olympiad-level math questions that are more abstract and complex. Here, $\mathcal{B}_{only\_math}$ has a slight edge (+1.8%) in the math category, which aligns with its domain-specific training. However, $\mathcal{B}{gpr\uparrow}$ demonstrates stronger performance in symbolic and language-heavy domains, showing +13.06% improvement in Law and +9.88% in English. Averaged across all non-math reasoning categories, $\mathcal{B}_{gpr\uparrow}$ achieves a +8.6% relative gain over $\mathcal{B}_{only\_math}$, reinforcing its advantage in general-purpose and real-world reasoning tasks.
Figure 7: Sub-category Accuracy Comparison across supergpqa. The $\mathcal{B}_{gpr\uparrow}$ blend consistently outperforms $\mathcal{B}_{only\_math}$ in a wide range of non-math reasoning categories except the science category which consists of fields like mathematics, physics, astronomy, chemistry etc.—highlighting the generalizability and versatility of multi-domain training.
A similar trend is observed in the supergpqa sub-category analysis shown in Figure 7. $\mathcal{B}_{gpr\uparrow}$ significantly outperforms $\mathcal{B}_{only\_math}$ across nearly all categories—especially in engineering, agronomy, economics, education, law, and philosophy. The only exception is the “Science” category, which includes math-heavy disciplines like physics, chemistry, and astronomy, where both blends perform comparably. This further highlights that multi-domain training with $\mathcal{B}_{gpr\uparrow}$ enhances reasoning across a broad spectrum of fields, achieving strong generalization even in real-world, professional domains that fall outside traditional math tasks.